OpenStudy (nuccioreggie):

58=12+2× Subtract 12 from both sides

OpenStudy (anonymous):

what do you get when you subtract 12 from both sides

OpenStudy (nuccioreggie):

Idk thats why I'm. Asking somone can they do it

OpenStudy (anonymous):

58-12

OpenStudy (nuccioreggie):

46

OpenStudy (anonymous):

12-12

OpenStudy (nuccioreggie):

0

OpenStudy (nuccioreggie):

0 @10115658

OpenStudy (anonymous):

\[58=12~+~2x\]\[58 - 12 = 12 - 12 + 2x\]\[46 = 2x\]\[\frac{46}{2} = \frac{2x}{2}\]